OpenStudy (anonymous):

f(x) = x-3 g(x) = square root x+5 what does (g o f ) (2) = ??

OpenStudy (anonymous):

find \[f(2)\] the find \[g(f(2))\]

OpenStudy (anonymous):

\[f(2)=2-3=-1\] \[g(-1)=\sqrt{-1+5}=\sqrt{4}=2\] fini